The fibonacci sequence is defined as x0 = 0, x1 = 1 and xn+2 = xn +xn+1 , for all non negative integers n prove that, xm = xr+1xm-r + xrxm-r for all integers m ≥ 1 and 0 ≤ r ≤ m-1 and xd divides xkd for all integers k and d
1
Expert's answer
2020-05-07T19:22:21-0400

1.


"x_m=x_{r+1}x_{m-r}+x_rx_{m-r}"

Prove by the induction


"x_m=x_{m-r}x_{r+2} \ \ \ \ \ \ \ \ \ \ \ \ (*)"

Since "m\geq1" the base of induction is:

"m=1=>r=0: x_1=x_{1-0}x_{0+2}=1\cdot1=1"

Hence, we proved the base of induction.

Now, we will assume, that the statement (*) is true for every "m<k." Let's prove for "m=k"

Prove by the induction (by r)

"r=0"


"x_k=x_{k-0}x_{0+2} =x_kx_2=x_k"

The induction assumption is that for every "r," such that: "0\leq r<r_0\leq k-1," the statement (*) is true. Let's prove it for "r=r_0." We should check the following:


"x_k=x_{k-r_0}x_{r_0+2}=x_{k-r_0}x_{r_0+1}+x_{k-r_0}x_{r_0}"

We can apply the induction assumption to both of the summands in the following way:

For the first summand:


"r_0+1=(r_0-1)+2, k-r_0=(k-1)-(r_0-1)"

We obtain the following(by using the (*)): 


"x_{k-r_0}x_{r_0+1}=x_{k-1}"

By applying the induction assumption and (*) to the second summand we obtain: 


"x_{k-r_0}x_{r_0}=x_{k-2}"

Thus we should check if:


"x_k=x_{k-1}+x_{k-2}"

This is true.


"x_m=x_{m-r}x_{r+2}=x_{m-r}(x_{r+1}+x_r)=x_{m-r}x_{r+1}+x_{m-r}x_r"

We prove


"x_m=x_{r+1}x_{m-r}+x_rx_{m-r}"

2.

"x_k|x_{kd}," where "k,d" are some integers.

Let's use the statement:


"x_m=x_{m-r}x_{r+2}""x_{kd}=x_{kd-d}x_{d+2}"

By applying this fact "(k-1)" times, we obtain the following: 


"x_{kd} = x_{d} x_{d+2}^{k-1}."

Thus "(k\geq1): x_k|x_{kd}," where "k,d" are some integers.